NOAA SATOPS Morning Report: Wednesday, January 16, 2013


SATOPS Morning Report: Wednesday, January 16, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 016

GOES
GOES-12 DOY015 image frame breaks due to BER at WCDA 12:58z,14:27z, 14:38z

GOES-12 1628Z Image degraded and 16:31z Sounding degraded due to an antenna tracking issue.

GOES-15 DOY015 1615Z PACUS image lost due to an antenna drive problem.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
OPS51 DOY015 14:35z Cat1 data loss of 6053 s/c seconds due to a LION ground station problem.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
ESPC reports that the 1615 UTC GOES-15 West data was lost due to an antenna problem at Wallops

GINI AWIPS West Images Lost; 1610 UTC West CONUS data was short so the product was not processed and the 1615 UTC data was not available due to the west antenna problem at Wallops and the 1630 UTC was skipped due to delays in attempting to process the 1615 UTC data.

DMSP/ SSMI F-15 rev 31983 was not received at ESPC

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.